OpenStudy (anonymous):

please help for medal The amount of money an electrician charges when visiting a home to make a repair is shown by the linear equation c = 110 + 85h, where c is the amount the electrician charges and h is the number of hours spent at the home. How much does the electrician charge when h = 3.5?

OpenStudy (anonymous):

@*Shadow*

OpenStudy (anonymous):

@CroKing

OpenStudy (anonymous):

@Jordan8901 @gerryliyana

OpenStudy (anonymous):

c = 110 + 85h c = 110 + 85(3.5) c = 110 + 297.5

OpenStudy (anonymous):

c = 407.5

OpenStudy (anonymous):

the amount the electrician charges 407.5

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

the amount the electrician charges 407.5 so you have got to be careful with how power you are useing now in your house